4x - y=9 is this ordered pair a solution of the given system of equations.

Found 2 solutions by rfer, ReadingBoosters:
Answer by rfer(16322)

What ordered pair?

Answer by ReadingBoosters(3246)

You didn't include the ordered pair.
...
Take the x value and input it into the equation and solve for y.
If y is the same as the one in the ordered pair, then that ordered pair is a solution.
...
For example, using the ordered pair (2,5)
4(2)-y=9
8-y=9
8-9=y
-1=y
Therefore, (2,5) is NOT a solution for that equation but (2,-1) is a solution.
...
Feel free to email me if you have more questions. homeworkhelpers@readingboosters.com
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.